A wooden disc of 2 m radius and 120 kg mass rotates through a vertical axis at the

centre of the disc. Later a 100 kg bag containing sand is released vertically onto the disc

at a point near the side of the disc. (o = 3 rad s-1)


Calculate the


(a) moment of inertia of the disc. [3 m]

(b) angular velocity after the sand bag is released (assume that the sand bag is a

particle).


1
Expert's answer
2021-10-06T08:13:41-0400

(a) The moment of inertia of the empty disk:


"I=\frac12 MR^2=240\text{ kg m}^2."

(b) Apply angular momentum conservation:


"L_i=L_f,\\
I\omega_0=(I+I_b)\omega,\\\space\\
\omega=\omega_0\frac{I}{I+I_b}=\omega_0\frac{I}{I+mR^2}=1.125\text{ rad/s}."